# Intervention Efficiency and Perturbation Validation Framework: Capacity-Aware and Robust Clinical Model Selection under the Rashomon Effect

### v6

@@ 1 Introduction
Small-scale clinical prediction modeling often faces challenges of limited sample size, class imbalance, and high dimensionality, with few established predictors. Data scarcity arises from high collection costs, ethical constraints, and follow-up difficulties (coors2017ethical). Class imbalance is common since adverse events affect only a small fraction of patients, leading to misleading results if conventional metrics like accuracy are used (davis2021accuracy; he2009learning). In such contexts, clinicians value precision and recall more than overall accuracy, especially when only a limited number of interventions can be performed (martin2025use; li2025mortality; gurudevan2024lung).
Because improving precision often reduces recall, an explicit trade-off is essential under constrained resources. The F1 score, as the harmonic mean of precision and recall, ignores true negatives and can favor suboptimal models when this trade-off is unclear (he2009learning). Threshold-independent metrics mitigate this issue but have drawbacks: AUC-ROC can overestimate performance on imbalanced data, while AUC-PR, though better for rare events, remains sensitive to prevalence and difficult to interpret clinically (saito2015precision; richardson2024receiver). Metrics such as balanced accuracy, Cohen’s kappa, and MCC have been proposed to address imbalance, with MCC noted for robustness (chicco2020advantages), yet all may yield ambiguous results requiring manual interpretation. Evaluating many models across several criteria is also computationally costly, motivating a measure that captures the precision–recall trade-off while reflecting capacity constraints.
High dimensionality further increases overfitting and obscures relevant predictors (berisha2021digital; bommert2022benchmark), motivating feature selection or dimensionality reduction (jolliffe2016principal; jia2022feature). Still, models like logistic regression, SVMs, k-nearest neighbors, and random forests can achieve similar accuracies while relying on distinct feature sets (lukaszuk2024stability). This phenomenon, called the Rashomon Effect (breiman2001statistical; rudin2024amazing; xin2022exploring), reveals that many “equally good” but mutually-exclusive models may coexist. This multiplicity usually complicated interpretation and clinical trust.
Model evaluation is also hindered by instability across data splits: small datasets amplify variance, and similar scores can occur by chance (cawley2010over). Traditional validation provides point estimates but neglects robustness, though model performance is known to fluctuate under realistic perturbations (varoquaux2018cross; paschali2018generalizability; arleto2023robust). Robustness to data shifts is now recognized as a core requirement for safe deployment in healthcare (paschali2018generalizability; pfohl2019benchmarking; heinrich2022robustness).
@@ 2 Related Work
Small, imbalanced datasets create high variance, overfitting, and instability in model and feature selection (cawley2010overfitting; zhang2023model; xin2022exploring), making it difficult to select one reliable model among many, particularly when the Rashomon Effect appears and causes multiplicity in models. In low-sample settings, small data perturbations can cause drastic changes in model structure and predictions, especially for unstable learners such as decision trees or stepwise regression (breiman1996bagging; breiman1996heuristics; morozova2015comparison; meinshausen2010stability; dietterich2000ensemble). Under class imbalance, models may favor majority classes or overfit rare ones (chawla2002smote; he2009learning; buda2018systematic; vandenGoorbergh2022harm).
Cross-validation (CV) is widely used to estimate model performance but performs unreliably on small, high-dimensional data. Specifically, CV-based selection often yields optimistic estimates and unstable rankings across splits (varma2006bias; cawley2010overfitting; arlot2010survey; varoquaux2018cross; zhang2021cross). Nested CV can reduce bias but remains computationally expensive and unstable when samples are few (wainer2018nested). Moreover, conventional CV averages over models rather than assessing individual consistency. Likewise, although ensemble and bagging methods reduce variance and improve generalization (breiman1996bagging; breiman2001random), they typically produce complex, less interpretable systems prone to overfitting (dietterich2000ensemble; buhlmann2002analyzing; opitz1999popular). In contrast, clinical use prioritizes transparent, single-model predictions for interpretability and deployment (dong2020survey; valente2021improving; zadeh2023comparative).
@@ 2.2 Perturbation-Based Validation in Small, Imbalanced Clinical Datasets
Robustness evaluation often involves injecting random, non-adversarial noise. Classic studies added Gaussian noise to features to assess performance degradation (egmont1997robustness; shafiee2019impact; parmar2015radiomic; yang2020ecg; pfohl2021impact; kulkarni2021gaussnoise). Recent gene–expression experiments found that logistic regression retained the most stable feature sets, while Random Forests were least stable, despite comparable accuracy (lukaszuk2024stability). This shows that conventional metrics alone can overlook robustness to noise.
For interpretable models, robustness work is emerging. For example, loss-function redesign improves single-tree resilience to label noise (sztukiewicz2024tree), yet trees in small-sample settings can still memorize noisy outliers, and ensemble averaging may leave feature importance unstable (soloff2024bagging).
Perturbation-based validation explicitly tests stability against noise. Label-noise injection, as in Mutation Validation (MV) (zhang2023model), flips a subset of labels to measure performance drop. Feature noise assesses covariate sensitivity (egmont1997robustness; kulkarni2021gaussnoise). Resampling methods, such as bootstrap, repeated k-fold CV and bagging, perturb sample composition to produce performance distributions rather than single estimates (breiman1996bagging; breiman2001random; soloff2024bagging).
@@
We also considered label perturbation during validation. Prior work shows that robust models should withstand limited label noise in training (zhang2023model). However, introducing label noise in validation can distort comparisons. Flipped labels disproportionately penalize stronger models (which previously predicted correctly) while leaving weaker ones largely unaffected, compressing or inverting true performance gaps. In highly imbalanced datasets, even a few flipped labels can drastically alter metrics like precision and recall. Therefore, PVF excludes label perturbation in validation, despite its popularity for robustness testing during training.
@@ 4.2 General Experimental Setup
After selection, the chosen model is either assessed based on ground truth information or when it is not available, as an approximation of it, on a large external test set using the same metric as in validation. To improve reliability and reduce variance from any single split, each configuration is repeated extensively. Model classes and other dataset-specific design choices are kept simple and interpretable to mimic clinical use (stiglic2020interpretability), and are described in subsequent subsections.
@@ 4.3 Overview of Experimental Results
In this subsection, similar to the synthetic dataset experiment, we also present the experimental results in the form of a comparison between the performance of PVF under $\sigma$ (defined as in Section 3.2) tuning and that of the traditional method, respectively for the cervical cancer (wolberg1993breast) and breast cancer (fernandes2017cervical) datasets. We refer the readers to Appendix D for the detailed setups of this part of experiments, including exactly how we calculate the performance gaps between PVF and the traditional method with repetitive experiments.
@@ C.1 Synthetic Dataset Experimental Setup
Clinical settings prioritize interpretability, often favoring simpler models like logistic regression or decision trees (abdullah2021review; rudin2019stop; Christodoulou2019). In practice, one can even decide the complexity of the model based on heuristic judgement on overfitting and underfitting. In this case, choosing the best model often reduces to choosing the right subset of features. This equivalence is clearest for bounded-depth decision trees or sparse logistic regression, whose predictive power hinges directly on the selected variables (breiman2001statistical; chen2022tree; ning2022interpretable; tan2022figs; nguyen2023interpretable; suttaket2024rational). Following these ideas, we use logistic regression models trained on pairs of features only. This yields ten pairs and thus ten model candidates, one per distinct pair among the five features. By the Rashomon Effect, there might be more than one pair of features that achieve optimal or almost-optimal performances. For example, although the truly informative pair is $(x_{1},x_{2})$, due to a small number of samples, $(x_{1},x_{3})$ might accidentally performs equally well or even better than $(x_{1},x_{2})$. Our design deliberately creates the possibility of encountering this situation under the Rashomon Effect, and aligns feature selection with model selection: the best model should be the one trained on the truly informative pair $(x_{1},x_{2})$. It creates a clear, ground-truth notion of “correct selection” while keeping the interpretability of the selected model.
@@ Appendix D Detailed Setups of Real Data Experiments
We evaluate on two public clinical datasets to ground the study in realistic settings. The Breast Cancer Wisconsin (Diagnostic) dataset (wolberg1993breast) has 569 samples, 212 positives, and no missing values; no preprocessing beyond standard scaling is needed. The Cervical Cancer (Risk Factors) dataset (fernandes2017cervical) has 808 samples and substantial missing values. We discard features with more than 50% missing values and impute the remaining entries with $k$-nearest neighbors ($k=5$). This yields 668 usable samples with 34 features. Among the four diagnostic targets, we use the Schiller test outcome because its higher prevalence provides adequate positive cases for stable validation.

### v7

@@ 1 Introduction
Small-scale clinical prediction modeling often faces challenges of limited sample size, class imbalance, and high dimensionality, with few established predictors. Data scarcity arises from high collection costs, ethical constraints, and follow-up difficulties (Coors et al. 2017). Class imbalance is common since adverse events affect only a small fraction of patients, leading to misleading results if conventional metrics like accuracy are used (Davis and Maiden 2021; He and Garcia 2009). In such contexts, clinicians value precision and recall more than overall accuracy, especially when only a limited number of interventions can be performed (Martin et al. 2025; Li et al. 2025; Gao et al. 2021).
Because improving precision often reduces recall, an explicit trade-off is essential under constrained resources. The F1 score, as the harmonic mean of precision and recall, ignores true negatives and can favor suboptimal models when this trade-off is unclear (He and Garcia 2009). Threshold-independent metrics mitigate this issue but have drawbacks: AUC-ROC can overestimate performance on imbalanced data, while AUC-PR, though better for rare events, remains sensitive to prevalence and difficult to interpret clinically (Saito and Rehmsmeier 2015; Richardson et al. 2024). Metrics such as balanced accuracy, Cohen’s kappa, and MCC have been proposed to address imbalance, with MCC noted for robustness (Chicco and Jurman 2020), yet all may yield ambiguous results requiring manual interpretation. Evaluating many models across several criteria is also computationally costly, motivating a measure that captures the precision–recall trade-off while reflecting capacity constraints.
High dimensionality further increases overfitting and obscures relevant predictors (Berisha et al. 2021; Bommert et al. 2022), motivating feature selection or dimensionality reduction (Jolliffe and Cadima 2016; Jia et al. 2022). Still, models like logistic regression, SVMs, k-nearest neighbors, and random forests can achieve similar accuracies while relying on distinct feature sets (Łukaszuk and Krawczuk 2024). This phenomenon, called the Rashomon Effect (Breiman 2001b; Rudin et al. 2024; Xin et al. 2022), reveals that many “equally good” but mutually-exclusive models may coexist. This multiplicity usually complicated interpretation and clinical trust.
Model evaluation is also hindered by instability across data splits: small datasets amplify variance, and similar scores can occur by chance (Cawley and Talbot 2010a). Traditional validation provides point estimates but neglects robustness, though model performance is known to fluctuate under realistic perturbations (Varoquaux 2018; Paschali et al. 2018; Arleto et al. 2023). Robustness to data shifts is now recognized as a core requirement for safe deployment in healthcare (Paschali et al. 2018; Pfohl et al. 2019; Heinrich et al. 2022).
@@ 2 Related Work
Small, imbalanced datasets create high variance, overfitting, and instability in model and feature selection (Cawley and Talbot 2010b; Zhang et al. 2023; Xin et al. 2022), making it difficult to select one reliable model among many, particularly when the Rashomon Effect appears and causes multiplicity in models. In low-sample settings, small data perturbations can cause drastic changes in model structure and predictions, especially for unstable learners such as decision trees or stepwise regression (Breiman 1996a, b; Morozova et al. 2015; Meinshausen and Bühlmann 2010; Dietterich 2000). Under class imbalance, models may favor majority classes or overfit rare ones (Chawla et al. 2002; He and Garcia 2009; Buda et al. 2018; van den Goorbergh et al. 2022).
Cross-validation (CV) is widely used to estimate model performance but performs unreliably on small, high-dimensional data. Specifically, CV-based selection often yields optimistic estimates and unstable rankings across splits (Varma and Simon 2006; Cawley and Talbot 2010b; Arlot and Celisse 2010; Varoquaux 2018; Zhang et al. 2021). Nested CV can reduce bias but remains computationally expensive and unstable when samples are few (Wainer and Cawley 2021). Moreover, conventional CV averages over models rather than assessing individual consistency. Likewise, although ensemble and bagging methods reduce variance and improve generalization (Breiman 1996a, 2001a), they typically produce complex, less interpretable systems prone to overfitting (Dietterich 2000; Bühlmann and Yu 2002; Opitz and Maclin 1999). In contrast, clinical use prioritizes transparent, single-model predictions for interpretability and deployment (Dong et al. 2020; Valente et al. 2021; Zadeh et al. 2023).
@@ 2.2 Perturbation-Based Validation in Small, Imbalanced Clinical Datasets
Robustness evaluation often involves injecting random, non-adversarial noise. Classic studies added Gaussian noise to features to assess performance degradation (Egmont-Petersen et al. 1997; Shafiee et al. 2019; Parmar et al. 2015; Yang et al. 2020; Pfohl et al. 2021; Kulkarni et al. 2021). Recent gene–expression experiments found that logistic regression retained the most stable feature sets, while Random Forests were least stable, despite comparable accuracy (Łukaszuk and Krawczuk 2024). This shows that conventional metrics alone can overlook robustness to noise.
For interpretable models, robustness work is emerging. For example, loss-function redesign improves single-tree resilience to label noise (Sztukiewicz et al. 2024), yet trees in small-sample settings can still memorize noisy outliers, and ensemble averaging may leave feature importance unstable (Soloff et al. 2024).
Perturbation-based validation explicitly tests stability against noise. Label-noise injection, as in Mutation Validation (MV) (Zhang et al. 2023), flips a subset of labels to measure performance drop. Feature noise assesses covariate sensitivity (Egmont-Petersen et al. 1997; Kulkarni et al. 2021). Resampling methods, such as bootstrap, repeated k-fold CV and bagging, perturb sample composition to produce performance distributions rather than single estimates (Breiman 1996a, 2001a; Soloff et al. 2024).
@@
We also considered label perturbation during validation. Prior work shows that robust models should withstand limited label noise in training (Zhang et al. 2023). However, introducing label noise in validation can distort comparisons. Flipped labels disproportionately penalize stronger models (which previously predicted correctly) while leaving weaker ones largely unaffected, compressing or inverting true performance gaps. In highly imbalanced datasets, even a few flipped labels can drastically alter metrics like precision and recall. Therefore, PVF excludes label perturbation in validation, despite its popularity for robustness testing during training.
@@ 4.2 General Experimental Setup
After selection, the chosen model is either assessed based on ground truth information or when it is not available, as an approximation of it, on a large external test set using the same metric as in validation. To improve reliability and reduce variance from any single split, each configuration is repeated extensively. Model classes and other dataset-specific design choices are kept simple and interpretable to mimic clinical use (Stiglic et al. 2020), and are described in subsequent subsections.
@@ 4.3 Overview of Experimental Results
In this subsection, similar to the synthetic dataset experiment, we also present the experimental results in the form of a comparison between the performance of PVF under $\sigma$ (defined as in Section 3.2) tuning and that of the traditional method, respectively for the cervical cancer (Wolberg et al. 1993) and breast cancer (Fernandes et al. 2017) datasets. We refer the readers to Appendix D for the detailed setups of this part of experiments, including exactly how we calculate the performance gaps between PVF and the traditional method with repetitive experiments.
@@ C.1 Synthetic Dataset Experimental Setup
Clinical settings prioritize interpretability, often favoring simpler models like logistic regression or decision trees (Abdullah et al. 2021; Rudin 2019; Christodoulou et al. 2019). In practice, one can even decide the complexity of the model based on heuristic judgement on overfitting and underfitting. In this case, choosing the best model often reduces to choosing the right subset of features. This equivalence is clearest for bounded-depth decision trees or sparse logistic regression, whose predictive power hinges directly on the selected variables (Breiman 2001b; Chen et al. 2022; Ning et al. 2022; Tan et al. 2022; Nguyen et al. 2023; Suttaket et al. 2024). Following these ideas, we use logistic regression models trained on pairs of features only. This yields ten pairs and thus ten model candidates, one per distinct pair among the five features. By the Rashomon Effect, there might be more than one pair of features that achieve optimal or almost-optimal performances. For example, although the truly informative pair is $(x_{1},x_{2})$, due to a small number of samples, $(x_{1},x_{3})$ might accidentally performs equally well or even better than $(x_{1},x_{2})$. Our design deliberately creates the possibility of encountering this situation under the Rashomon Effect, and aligns feature selection with model selection: the best model should be the one trained on the truly informative pair $(x_{1},x_{2})$. It creates a clear, ground-truth notion of “correct selection” while keeping the interpretability of the selected model.
@@ Appendix D Detailed Setups of Real Data Experiments
We evaluate on two public clinical datasets to ground the study in realistic settings. The Breast Cancer Wisconsin (Diagnostic) dataset (Wolberg et al. 1993) has 569 samples, 212 positives, and no missing values; no preprocessing beyond standard scaling is needed. The Cervical Cancer (Risk Factors) dataset (Fernandes et al. 2017) has 808 samples and substantial missing values. We discard features with more than 50% missing values and impute the remaining entries with $k$-nearest neighbors ($k=5$). This yields 668 usable samples with 34 features. Among the four diagnostic targets, we use the Schiller test outcome because its higher prevalence provides adequate positive cases for stable validation.